OpenStudy (anonymous):

how do you solve proportions like 1/3=x/12

hero (hero):

Cross multply

hero (hero):

If \[\frac{a}{b} = \frac{c}{d}\]Then \[a \times d = b \times c\]

OpenStudy (bibby):

Algebra. Cross multiplication doesn't make sense if you don't understand why it works \[\huge \frac{12 }{ 1 } *\frac{ 1 }{ 3 } = \frac{ x }{ \cancel{12} }*\frac{ \cancel{12} }{ 1 }\]